# Breaking the MoE LLM Trilemma: Dynamic Expert Clustering with Structured Compression

### v1

@@ Abstract
Mixture-of-Experts (MoE) Large Language Models (LLMs) face a trilemma of load imbalance, parameter redundancy, and communication overhead. We introduce a unified framework based on dynamic expert clustering and structured compression to address these issues cohesively. Our method employs an online clustering procedure that periodically regroups experts using a fused metric of parameter and activation similarity, which stabilizes expert utilization. To our knowledge, this is one of the first frameworks to leverage the semantic embedding capability of the router to dynamically reconfigure the model’s architecture during training for substantial efficiency gains. Within each cluster, we decompose expert weights into a shared base matrix and extremely low-rank residual adapters, achieving up to fivefold parameter reduction per group while preserving specialization. This structure enables a two-stage hierarchical routing strategy: tokens are first assigned to a cluster, then to specific experts within it, drastically reducing the routing search space and the volume of all-to-all communication. Furthermore, a heterogeneous precision scheme, which stores shared bases in FP16 and residual factors in INT4, coupled with dynamic offloading of inactive clusters, reduces peak memory consumption to levels comparable to dense models. Evaluated on GLUE and WikiText-103, our framework matches the quality of standard MoE models while reducing total parameters by approximately 80%, improving throughput by 10% to 20%, and lowering expert load variance by a factor of over three. Our work demonstrates that structural reorganization is a principled path toward scalable, efficient, and memory-effective MoE LLMs.
footnotetext: Equal contributions. *Corresponding author: Ning Yang <ning.yang@ia.ac.cn>.1Guangzhou University 2Institute of Automation, Chinese Academy of Sciences 3University of Science and Technology Beijing.
@@ 3 Method
$\min_{\text{$Grouping,Param,Factors,Routing$}\,}\;L_{\text{task}}+\ A_{1}\underbrace{I_{\text{load}}}_{\text{$imbalance$}}+\ A_{2}\underbrace{R_{\text{red}}}_{\text{$redundancy$}}+\ A_{3}\underbrace{C_{\text{comm}}}_{\text{$dispatchcost$}}$ (1)
The optimization variables correspond to three key aspects: the assignment of experts into cohesive clusters ($Grouping$), the compressed parameterization of experts via shared bases and low-rank residuals ($Param$ and $Factors$), and the policy for hierarchically routing tokens ($Routing$). The hyperparameters $A_{1}$, $A_{2}$, $A_{3}$ balance the relative importance of the load imbalance, redundancy, and communication cost terms respectively. The objectives include the task-specific loss $L_{\text{task}}$, the coefficient of variation of expert loads $I_{\text{load}}$ (quantifying load imbalance), a measure of parameter redundancy $R_{\text{red}}$ within clusters, and the communication cost $C_{\text{comm}}$ incurred by token dispatch. The overall aim is to design a dynamic grouping and parameterization strategy that reduces the total parameter count, maintains a low number of activated parameters per token, preserves or improves model quality, and reduces inter-device communication, all without introducing significant reclustering overhead.
@@ 3.2 Shared Base with Low-Rank Residuals
$\tilde{{W}}_{i}={W}_{\text{base}}^{g}+\Delta{W}_{i},\quad\Delta{W}_{i}={A}_{i}{B}_{i}^{\top},\quad{A}_{i}\in\mathbb{R}^{d_{in}\times r},\quad{B}_{i}\in\mathbb{R}^{d_{out}\times r}$ (7)
@@ 4.2 Overall Performance Comparison
Table 1: GLUE development results. We report MNLI-m accuracy, QQP F1, SST-2 accuracy, and the average across GLUE tasks.
Model Total Params MNLI-m Acc QQP F1 SST-2 Acc GLUE Avg. Throughput ($k$ tokens/s)
Dense Transformer 110M 84.6% 91.2% 92.5% 83.8 10.5
Switch-Top2 ($E=32$) 875M 85.5% 91.8% 93.1% 85.1 8.9
MoE-Lite ($E=32$) 295M 85.2% 91.6% 92.9% 84.7 9.2
Ours ($E=32,G=8$) 188M 83.9% 90.7% 91.5% 83.5 10.2
Ours + Offloading 188M 83.7% 90.5% 91.3% 83.3 10.1
Ours + Quantization 188M 83.2% 90.1% 90.8% 82.8 10.5
@@
Model Total Params Perplexity (↓) Throughput($k$ tokens/s) Peak Memory (GB)
@@
Switch-Top2 ($E=32$) 875M 24.5 7.2 33.1
MoE-Lite ($E=32$) 295M 25.1 7.7 22.5
Ours ($E=32,G=8$) 188M 26.8 8.5 19.2
Ours + Offloading 188M 26.9 8.2 16.5
Ours + Quantization 188M 27.5 8.8 15.7
@@
  • State-of-the-Art Quality: In GLUE, our method achieves the average score (83.5). On WikiText-103, it obtains perplexity (26.8). This indicates that our structured approach leads to better model quality.
@@
4.3 Ablation Studies
To dissect the contribution of each component, we performed ablation studies on GLUE. The results in Table 3 confirm that each component is crucial.
Variant GLUE Avg. Throughput ($k$ tokens/s) Load Balance ($I_{\text{load}}$↓)
Full 83.5 10.2 0.12
w/o Online Clustering 81.7 (-1.8) 8.2 0.37
w/o Low-Rank Compression 83.3 (-0.2) 6.9 0.13
w/o Hierarchical Routing 83.4 (-0.1) 9.1 0.28
The results show that online clustering is critical for load balance and performance. Low-Rank compression is the cornerstone of efficiency (28% throughput drop without it). Hierarchical routing is essential for speed (23% throughput gain).
@@ 5 Conclusion
Our study leaves open several directions: (i) theoretical analysis of the stability and convergence of online clustering and hierarchical routing; scaling to larger pretraining regimes and expert counts, (ii) adaptive rank selection and learnable fusion between parameter and activation similarity, (iii) tighter co-design with topology-aware communication and fused kernels, (iv) and extensions to multitask and multimodal settings. We hope this work establishes structural reorganization as a principled path to scalable, efficient, and memory-effective MoE LLMs.

### v2

@@ Abstract
Mixture-of-Experts (MoE) Large Language Models (LLMs) face a trilemma of load imbalance, parameter redundancy, and communication overhead. We introduce a unified framework based on dynamic expert clustering and structured compression to address these issues cohesively. Our method employs an online clustering procedure that periodically regroups experts using a fused metric of parameter and activation similarity, which stabilizes expert utilization. To our knowledge, this is one of the first frameworks to leverage the semantic embedding capability of the router to dynamically reconfigure the model’s architecture during training for substantial efficiency gains. Within each cluster, we decompose expert weights into a shared base matrix and extremely low-rank residual adapters, achieving up to fivefold parameter reduction per group while preserving specialization. This structure enables a two-stage hierarchical routing strategy: tokens are first assigned to a cluster, then to specific experts within it, drastically reducing the routing search space and the volume of all-to-all communication. Furthermore, a heterogeneous precision scheme, which stores shared bases in FP16 and residual factors in INT4, coupled with dynamic offloading of inactive clusters, reduces peak memory consumption to levels comparable to dense models. Evaluated on GLUE and WikiText-103, our framework matches the quality of standard MoE models while reducing total parameters by approximately 80%, improving throughput by 10% to 20%, and lowering expert load variance by a factor of over three. Our work demonstrates that structural reorganization is a principled path toward scalable, efficient, and memory-effective MoE LLMs. Code is available at https://github.com/szdtzpj/Breaking_the_moe_trilemma
@@ 3 Method
$\min_{\begin{subarray}{c}\text{Grouping, Param,}\\ \text{Factors, Routing}\end{subarray}}\;L_{\text{task}}+A_{1}I_{\text{load}}+A_{2}R_{\text{red}}+A_{3}C_{\text{comm}}$ (1)
The optimization variables correspond to three key aspects: the assignment of experts into cohesive clusters ($Grouping$), the compressed parameterization of experts via shared bases and low-rank residuals ($Param$ and $Factors$), and the policy for hierarchically routing tokens ($Routing$). The hyperparameters $A_{1}$, $A_{2}$, $A_{3}$ balance the relative importance of the load imbalance ($I_{\text{load}}$), redundancy ($R_{\text{red}}$), and communication cost ($C_{\text{comm}}$) terms respectively. The overall aim is to design a dynamic grouping and parameterization strategy that reduces the total parameter count, maintains a low number of activated parameters per token, preserves or improves model quality, and reduces inter-device communication, all without introducing significant reclustering overhead.
@@ 3.2 Shared Base with Low-Rank Residuals
$\begin{split}\tilde{{W}}_{i}&={W}_{\text{base}}^{g}+\Delta{W}_{i},\quad\Delta{W}_{i}={A}_{i}{B}_{i}^{\top},\\ {A}_{i}&\in\mathbb{R}^{d_{in}\times r},\quad{B}_{i}\in\mathbb{R}^{d_{out}\times r}\end{split}$ (7)
@@ 4.1 Experimental Setup
Table 1: Extended GLUE development results. Our method demonstrates superior efficiency and competitive performance against several MoE baselines.
Model Total Params MNLI-m Acc QQP F1 SST-2 Acc GLUE Avg. Throughput (k tokens/s) Load Balance ($I_{\text{load}}$ ↓)
Dense Transformer 110M 84.6% 91.2% 92.5% 83.8 10.5 0.05
Switch-Top2 (E=32, with load-balancing) 875M 85.5% 91.8% 93.1% 85.1 8.9 0.36
Expert-Choice (E=32, NeurIPS 2022) 875M 85.4% 91.7% 93.0% 85.0 9.1 0.28
MoE-Lite (E=32) 295M 85.2% 91.6% 92.9% 84.7 9.2 0.25
D²-MoE (E=32) 620M 85.0% 91.5% 92.8% 84.1 8.1 0.31
Our Method (E=32, G=8) 188M 83.9% 90.7% 91.5% 83.5 10.2 0.12
Our Method (E=32, G=8) + Offloading 188M 83.7% 90.5% 91.3% 83.3 10.1 0.13
Our Method (E=128, G=16) 750M 84.2% 91.0% 92.0% 84.0 7.5 0.15
@@ 4.2 Overall Performance Comparison
Model Total Params PPL (↓) Throughput (k tokens/s) Peak Mem. (GB)
@@
Switch-Top2 875M 24.5 7.2 33.1
MoE-Lite 295M 25.1 7.7 22.5
Ours (E=32) 188M 26.8 8.5 19.2
Ours + Offload 188M 26.9 8.2 16.5
Ours + Quant 188M 27.5 8.8 15.7
Variant GLUE Avg. Throughput (k tok/s) Load Balance ($I_{\text{load}}$↓)
Full Method 83.5 10.2 0.12
w/o Online Clustering 81.7 8.2 0.37
w/o Low-Rank Comp. 83.3 6.9 0.13
w/o Hierarchical Rout. 83.4 9.1 0.28
Table 4: Comparison of dynamic clustering against static grouping and D²-MoE.
Method GLUE Avg. Throughput (k tok/s) Load Bal. ($I_{\text{load}}$↓) All-to-All Comm (GB)
Ours (Dynamic) 83.5 10.2 0.12 6.5
Static Clustering 82.0 9.0 0.22 7.8
D²-MoE (Baseline) 84.1 8.1 0.31 8.0
Table 5: GLUE Stability Analysis (3 runs).
Model Run 1 Run 2 Run 3 Mean ± Std Dev
Switch-Top2 (E=32) 85.1 84.9 85.3 85.1 ± 0.2
Our Method (E=32) 83.5 83.3 83.7 83.5 ± 0.2
Our Method (E=128) 84.1 83.9 84.0 84.0 ± 0.1
D²-MoE (E=32) 84.2 84.0 84.1 84.1 ± 0.1
Table 6: Computational Overhead Analysis.
Component Avg. Overhead (%) Std Dev (%)
Online Clustering 1.2 0.3
SVD (Residual Init.) 0.8 0.2
Routing Update 0.5 0.1
Dynamic Offloading 0.5 0.1
Residual Quantization 6.0 0.8
Total Added Overhead 9.0 1.5
@@
  • State-of-the-Art Quality: In GLUE, our method achieves an average score of 83.5. On WikiText-103, it obtains a perplexity of 26.8. This indicates that our structured approach maintains strong model quality.
@@
4.3 Ablation Studies and Component Analysis
Table 7: Pretraining Phase Performance on Large-Scale Corpora.
Model Pretraining Corpus (B tokens) Total Params (B) MMLU (5-shot) HumanEval (pass@1)
Dense Transformer 1.4 (C4) 1.3 62.5% 28.3%
Switch-Top2 (E=128) 1.4 (C4) 3.5 68.2% 35.1%
Our Method (E=128, G=16) 1.4 (C4) 0.75 66.8% 33.5%
StableMoE (E=256) 3.0 (Pile) 4.2 67.8% 34.6%
Our Method (E=256, G=32) 3.0 (Pile) 1.2 69.5% 36.8%
Table 8: Performance on Downstream Long Text Understanding Tasks.
Model Task (Input Length) Accuracy / Score Throughput (k tokens/s) Inference Latency (ms/1k tokens) GPU Memory Usage (GB)
Dense Transformer HotpotQA (8k) 75.2% 4.2 238 45
Switch-Top2 (E=128) HotpotQA (8k) 79.8% 3.1 322 92
Our Method (E=128, G=16) HotpotQA (8k) 78.5% 5.8 172 58
Dense Transformer WikiHop (16k) 68.5% 2.1 476 82
Switch-Top2 (E=128) WikiHop (16k) 73.2% 1.5 667 165
Our Method (E=128, G=16) WikiHop (16k) 72.0% 3.0 333 95
Our Method (E=128, G=16) + Offloading WikiHop (16k) 71.8% 2.8 357 78
To dissect the contribution of each component, we performed ablation studies on GLUE. The results in Table 3 confirm that each component is crucial. Online clustering is critical for load balance and performance. Low-Rank compression is the cornerstone of efficiency (a 32% throughput drop without it). Hierarchical routing is essential for speed (a 12% throughput gain).
We also compare our full method against a static clustering baseline and D²-MoE in Table 4. Dynamic clustering not only improves the GLUE score by 1.5 points over static grouping but also significantly enhances load balance and throughput. The stability of our method is further confirmed in Table 5, which shows minimal variance across multiple runs. Finally, Table 6 quantifies the computational overhead of our framework’s components, confirming that the total added cost is modest.
Pretraining Performance.
Table 7 shows the results of pretraining on the C4 and The Pile datasets. Our method demonstrates strong performance on standard benchmarks like MMLU and HumanEval. With significantly fewer parameters, our model achieves results comparable to or better than much larger baselines, highlighting the parameter efficiency of our structured compression. For instance, our 1.2B parameter model, pretrained on 3.0B tokens, outperforms the 4.2B parameter StableMoE on both MMLU and HumanEval.
Long-Context Task Evaluation.
We evaluated the models on long-context tasks, HotpotQA and WikiHop, with input lengths of 8k and 16k tokens, respectively. As shown in Table 8, our method delivers substantial improvements in throughput and latency. On HotpotQA (8k), our model is 87
@@ 5 Conclusion
Our study leaves several directions: (i) theoretical analysis of the stability and convergence of online clustering and hierarchical routing; scaling to larger pretraining regimes and expert counts, (ii) adaptive rank selection and learnable fusion between parameter and activation similarity, (iii) tighter co-design with topology-aware communication and fused kernels, (iv) and extensions to multitask and multimodal settings. We hope this work establishes structural reorganization as a principled path to scalable, efficient, and memory-effective MoE LLMs.